# Pinching Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

### v1

@@ Abstract
This article proposes a novel design for the Pinching Antenna Systems (PASS) and advocates simple yet efficient wireless communications over the ‘last meter’. First, the potential benefits of PASS are discussed by reviewing an existing prototype. Then, the fundamentals of PASS are introduced, including physical principles, signal models, and communication designs. In contrast to existing multi-antenna systems, PASS brings a novel concept termed Pinching Beamforming, which is achieved by dynamically adjusting the positions of PAs. Based on this concept, a couple of practical transmission architectures are proposed for employing PASS, namely non-multiplexing and multiplexing architectures. More particularly, 1) The non-multiplexing architecture is featured by simple baseband signal processing and relies only on the pinching beamforming; while 2) the multiplexing architecture provides enhanced signal manipulation capabilities with joint baseband and pinching beamforming, which is further divided into sub-connected, fully-connected, and phase-shifter-based fully-connected schemes. Furthermore, several emerging scenarios are put forward for integrating PASS into future wireless networks. As a further advance, by demonstrating a few numerical case studies, the significant performance gain of PASS is revealed compared to conventional multi-antenna systems. Finally, several research opportunities and open problems of PASS are highlighted.
@@ I Introduction
The perspective introduced by Shannon’s information theory established the foundation for the digital revolution and has profoundly shaped the development of modern communication systems [1]. In Shannon’s original framework, the channel connecting the transmitter and receiver was treated as a fixed entity, which is entirely dictated by environmental factors and beyond human control. However, advances in communication technology have challenged this assumption by enabling active reconfiguration of channel parameters to boost the channel throughput. A notable example is the deployment of multiple-input multiple-output (MIMO) systems, which reconfigures the channel by creating parallel point-to-point links to enhance communication efficiency [2].
Building on this momentum, recent advancements in intelligent and flexible antenna design have introduced enhanced approaches methods for channel reconfiguration. Technologies such as reconfigurable intelligent surfaces [3, 4], dynamic metasurface antennas [5], fluid antennas [6], and movable antennas [7] can dynamically adjust their positions, apertures, and electromagnetic properties, enabling precise control of wireless channel conditions. By optimizing these reconfigurable parameters, these novel antenna types and array architectures offer flexibility in controlling the effective end-to-end channel gain. This capability has demonstrated great potential for improving system performance and has attracted considerable research interest.
Despite their potential to enhance network throughput, existing flexible-antenna systems face certain limitations. Firstly, antenna reconfiguration is often confined to apertures spanning only a few wavelengths, which limits their ability to mitigate large-scale path loss. Secondly, in existing reconfigurable systems aforementioned, it is costly to add or remove antennas after establishing. To address these challenges, a more flexible yet efficient new design—pinching-antenna systems (PASS) come into researchers’ sight [8, 9].
@@ I-A The Road to Near-Wired: A Prototype of PASS
The first prototype of PASS was developed and tested by NTT DOCOMO in 2021 [8], as illustrated in Figure 1. A PASS comprises a type of specialized transmission line referred to as a “dielectric waveguide”, which works as a leaky-wave antenna by pinching small dielectric particles—known as separate dielectrics—at specific points along its length. These dielectrics are often attached to the tips of plastic clips or clothespins, forming what are called pinching antennas (PAs). By exploiting the ability of the waveguide to transmit or receive radio waves at any point pinched with a dielectric, PASS enable wireless communications at any location around the pinched point [8]. This ability allows PASS to swiftly establish a variety of communication areas tailored to specific locations and environmental conditions.
The waveguide can be pinched at any point along its length to provide coverage at the corresponding location, including multiple points along the line. Removing the clip immediately stops signal transmission at that specific location. This process is similar to adding or removing a clothespin from a clothesline, as illustrated in Figure 1, and offers exceptional flexibility in antenna deployment [8].
In contrast to conventional flexible-antenna systems, the waveguide’s length of PASS can be extended arbitrarily, which allows PAs to be positioned close to users to establish strong line-of-sight (LoS) links [9]. This adaptability effectively creates a near-wired connection and thus ensures favorable channel conditions, which shifts the focus of wireless communication networks from the “last mile” to the more manageable “last meter”. This unique capability is further encapsulated in the abbreviation “PASS”, which emphasizes PASS’s ability to provide the shortest and most efficient wireless connections between antennas and users to pass wireless signals.
@@ I-B Key Advantages for Employing PASS in Wireless Communications
  • Enhanced LoS Stability and “Last Meter” Communications: By extending the waveguide and enabling the flexible deployment of PAs, PASS effectively mitigate large-scale path loss and establish stable LoS communication environments. This capability facilitates “last meter” communications and expands network coverage, which is not achievable with other flexible-antenna technologies.
  • Flexible Pinching Beamforming Gain: PASS control signals by optimizing the locations and numbers of PAs, which enables highly flexible and scalable pinching beamforming. This capability allows precise management of signal strength and interference, effectively meeting stringent service requirements.
  • Utilization of Near-Field Benefits: The extended aperture of the waveguide places users within the near-field region of PASS, which enables the use of near-field electromagnetic properties, such as beam focusing, to enhance wireless transmission efficiency [10].
  • Cost-Effective and Scalable Implementation: PASS are simple to implement and economical to deploy, as they only require adding or removing dielectric materials at selected points along the waveguide [8], which makes PASS quite scalable choice for practical network design.
These advantages have fueled growing research interest in PASS. The performance of PASS has been analyzed in terms of its ergodic rate [9] and array gain [11]. Furthermore, various activation algorithms have been proposed to optimize the locations of PAs along a single waveguide; see [9] and relevant references. However, despite these initial efforts, the design of PASS for wireless communications remains in its early stages. This gap motivates us to provide a systematic introduction to PASS. In this article, we first discuss how PASS differ from conventional multi-antenna systems, focusing on the fundamental physical principles, signal model, and communication system design. We then propose several novel beamforming architectures to facilitate the design of pinching beamforming. This is followed by an exploration of promising application scenarios for PASS and an evaluation of their performance. Finally, we highlight the open challenges of PASS to inspire future research in this innovative area.
@@ II-A Physical Principle
Unlike conventional electronic antennas which rely on electronic devices to convert alternating electric current into radio waves, the physical principle of PAs is based on the phenomenon of electromagnetic coupling that arises when a separate dielectric is brought into close proximity with a dielectric waveguide [8]. This coupling enables the power exchange between the waveguide and the separate dielectric, as illustrated in the bottom-right of Fig. 1. Therefore, by pinching a separate dielectric onto the waveguide, the radio waves originally confined within the waveguide leak into the dielectric, where they are subsequently radiated into free space, thereby forming an “antenna”. Unlike conventional electronic antennas, PAs are entirely passive, simpler to manufacture and install, and offer greater position flexibility.
By modeling the pinching antenna as an open-ended directional coupler, the power exchange caused by the electromagnetic coupling can be characterized by the coupled-mode theory. Let $P_{\mathrm{guide}}$ and $P_{\mathrm{pinch}}$ denote the normalized power in the waveguide and the pinched separate dielectric, respectively. Their relationship is approximately given by [12]
@@
where $L$ is the coupling length shown in Fig. 1, $\kappa$ is the coupling coefficient, and $F\leq 1$ is the maximum coupling efficiency. The factors $\kappa$ and $F$ are influenced by various factors, including the effective refractive index and the transverse modes of both the waveguide and the separate dielectric. It can be observed from the above formulas that the power radiated from a PA can be controlled by adjusting the coupling length $L$. When the waveguide and the separate dielectric share the same effective refractive index, the coupling efficiency $F$ reaches its maximum value of $1$, making it possible to radiate full power from a single PA with a coupling length of $\pi/(2\kappa)$.
@@ II-B Signal Model
As shown in Fig. 1, the composite channel from the base station (BS) to the user in the PASS system consists of two components: waveguide propagation and free-space propagation. Compared to free-space propagation, waveguide propagation incurs almost negligible path loss, e.g., approximately 0.01-0.03 dB/m for circular copper waveguide at 15 GHz [13]. Consequently, when multiple PAs are deployed on a waveguide, the signal received at the user from the $n$-th PA through the line-of-sight channel can be expressed mathematically as:
@@
Based on the physical principle of PAs, the power model of PASS, signified by $P_{n}$, is significantly different from that of conventional multi-antenna systems. Specifically, in conventional antennas, the power of the signal emitted from each antenna is actively controlled by electronic devices such as radio-frequency (RF) chains and power splitters. In contrast, for PAs, the transmit power is passively determined by the coupling length with the waveguide. Furthermore, when multiple PAs are pinched along a waveguide, the power radiated by subsequent antennas is influenced by the power radiated by preceding ones, leading to a coupled power model. While the coupling length of a PA is not as easily adjustable as the electronic parameters of conventional antennas, it is still possible to design and manufacture PAs with specific coupling lengths to achieve a desired power model. In the following, we introduce two common power models for PAs.
  • Equal Power Model: In this model, each PA on the waveguide radiates an equal amount of power. This can be achieved by designing the PAs with varying coupling lengths, such that the earlier antennas radiate a smaller proportion of the power in the waveguide, while the later antennas radiate a larger proportion of the remaining power.
  • Proportional Power Model: In this model, each PA is designed with an identical coupling length, thereby radiating the same fixed proportion of the power left in the waveguide. As a result, the power radiated by each antenna decreases progressively along the waveguide, where the power radiated by later antennas is proportional to that of earlier ones by a fixed ratio.
These two power models offer different pros and cons. Compared to the proportional power model, where the later PAs may radiate only a small amount of power, the equal power model ensures that all PAs are effectively utilized. The equal power model also simplifies the performance analysis of the PASS system due to its simple mathematical formulation, which has contributed to its widespread adoption in the literature. In contrast, the proportional power model requires less stringent manufacturing specifications, as all PAs share the same design, thus reducing production complexity and costs.
@@ II-C Communication System Design
From a communication system design perspective, it is crucial to position the PAs at suitable locations to maximize communication performance. On the one hand, the PAs should be deployed close to the user to mitigate the significant signal attenuation caused by free-space path loss and blockage. On the other hand, when multiple PAs are deployed along the waveguide, they form an antenna array, enabling pinching beamforming to enhance signal strength and reduce interference. By fine-tuning the positions of PAs along the waveguide, the phase shifts of the signals, resulting from propagation within the waveguide and free space, can be adjusted, thereby optimizing the beamforming gain. Leveraging the flexibility and portability of PAs, their positions can be determined either actively by the BS or passively by the user, leading to the following two designs:
  • BS-centric Design: In this design, the positions of PAs are determined by the BS to optimize specific performance metrics of the overall PASS system, such as spectral efficiency and user fairness. In dynamic scenarios, the positions of PAs can be adjusted actively based on the locations of all users in two ways: either by moving them along a pre-installed track parallel to the waveguide for continuous activation, or by selectively activating pre-installed PAs on the waveguide for discrete activation.
  • User-centric Design: In this design, users can pinch or release portable PAs onto the waveguide as needed to enhance communication service in their local area. Whenever a PA is pinched or released by a user, the baseband processing strategy needs to be optimized and updated based on the new propagation environment to minimize negative impacts on other users.
@@ III-A Non-Multiplexing Architecture
A schematic illustration of the non-multiplexing architecture is shown in Fig. 2. In this architecture, the data streams to be transmitted are not multiplexed at the baseband, meaning that each waveguide is fed by a single data stream. In this setup, the primary role of the baseband processing is to manage the power allocation for each data stream to optimize transmission efficiency. The key benefit of this design is to reduces implementation complexity. Compared to the conventional multi-antenna systems, the unique capability of pinching beamforming, achieved by adjusting the positions of PAs, can be utilized to enhance communication performance. This non-multiplexing architecture is particularly well-suited for scenarios where each waveguide serves geographically isolated areas, minimizing the need for advanced interference management between waveguides.
@@ III-B Multiplexing Architecture
As shown at the bottom of Fig. 3, the multiplexing architecture multiplexes multiple data streams before feeding them into the waveguide, thus facilitating the joint transmit and pinching beamforming. This architecture can be seen as an extension of the conventional MIMO transmission architecture, where each electronic antenna is replaced by a waveguide. The key difference between the two architectures lies in the point of signal radiation. In conventional MIMO systems, signals are radiated directly from the electronic antennas into free space. In contrast, in the multiplexing architecture of PASS, signals are radiated at specific locations along the waveguide via PAs. Consequently, unlike conventional MIMO, which relies on a massive number of electronic antennas to achieve high spatial degrees of freedom (DoFs) [14], the multiplexing architecture of PASS significantly reduces the requirement for numerous waveguides because a large number of low-cost PAs can be easily installed on a single waveguide. Based on the above discussion, the multiplexing architecture for PASS can be designed in analogy to conventional MIMO architecture but with distinct differences in implementation and design focus. In the following, we introduce three basic multiplexing architectures, as shown in Fig. 3, namely sub-connected architecture, fully-connected architecture, and phase shifter (PS)-based fully-connected architecture.
  • Sub-connected Architecture: This is the simplest multiplexing architecture, where each waveguide is connected to a dedicated RF chain and carries signals that are multiplexed and precoded by baseband digital beamforming. While this architecture is analogous to the fully-digital architecture used in conventional MIMO systems, it is referred to as “sub-connected” because the signals from each RF chain are emitted exclusively from the PAs deployed on a single waveguide.
  • Fully-connected Architecture: In this architecture, the signals from each RF chain are fed into all waveguides with the aid of power splitters. The fully-connected architecture addresses two key limitations of the sub-connected design. First, it enables each RF chain to connect to all PAs across all waveguides, thereby maximizing the utilization of the overall spatial DoFs to enhance communication performance. Second, the number of RF chains no longer needs to match the number of waveguides. This flexibility allows for either reducing the number of RF chains to save energy or increasing it to further exploit the additional spatial DoFs provided by the massive PAs.
@@
IV ‘PASS’ Signals for 6G and Beyond

### v2

# Pinching-Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

@@ Abstract
This article proposes a novel design for the Pinching Antenna SyStems (PASS) and advocates simple yet efficient wireless communications over the “last meter”. First, the potential benefits of PASS are discussed in the paper by reviewing an existing prototype. Then, the fundamentals of PASS are introduced, including their physical principles, signal models, and communication designs. In contrast to existing multi-antenna systems, PASS brings a novel concept, termed Pinching Beamforming, which is achieved by dynamically adjusting the positions of pinching antennas (PAs). Based on this concept, a couple of practical transmission architectures are proposed for realizing the full potentia of PASS, namely non-multiplexing and multiplexing architectures. More particularly, 1) the non-multiplexing architecture is featured by simple baseband signal processing and relies on the pinching beamforming only; while 2) the multiplexing architecture provides enhanced signal processing capabilities with joint baseband and pinching beamforming, which can be further divided into sub-connected, fully-connected, and phase-shifter-based fully-connected schemes. Furthermore, several emerging scenarios are put forward for integrating PASS into future wireless networks. As a further advance, by demonstrating a few numerical case studies, the significant performance gain of PASS is revealed compared to conventional multi-antenna systems. Finally, several research opportunities and open problems of PASS are highlighted.
@@ I Introduction
The perspective introduced by Shannon’s information theory established the foundation for the digital revolution and has profoundly shaped the development of modern communication systems [1]. In Shannon’s original framework, the channel connecting the transmitter and receiver was treated as a fixed entity, i.e., a parameter entirely dictated by environmental factors and beyond human control. However, advances in communication technology have challenged this assumption by enabling active reconfiguration of channel parameters to boost the channel throughput. A notable example is the deployment of multiple-input multiple-output (MIMO) systems, which reconfigures the channel by creating separated parallel point-to-point links to enhance communication efficiency [2].
Building on this momentum, recent advancements in intelligent and flexible antenna designs have further introduced novel approaches for channel reconfiguration. Technologies, such as reconfigurable intelligent surfaces [3, 4], dynamic metasurface antennas [5], fluid antennas [6], and movable antennas [7], can dynamically adjust their positions, apertures, and electromagnetic properties, enabling precise control of wireless channel conditions. By optimizing their reconfigurable parameters, these novel antenna types and array architectures offer flexibility in controlling the effective end-to-end channel gain. This capability has demonstrated great potential for improving system performance and has attracted considerable research interests.
Despite their potentials to enhance network throughput, existing flexible-antenna systems face certain limitations. Firstly, those existing antenna reconfigurations are often confined to apertures spanning only a few wavelengths, which limits their ability to mitigate large-scale path loss. Secondly, in existing reconfigurable systems, it is costly, if not impossible, to change their antenna configuration, i.e., add or remove antennas once these systems are established. To address these challenges, a more flexible yet efficient new design—Pinching-Antenna SyStems (PASS) come into researchers’ sight [8, 9].
@@ I-A The Road to Near-Wired: A Prototype of PASS
The first prototype of PASS was developed and demonstrated by NTT DOCOMO in 2021 [8], as illustrated in Figure 1. A PASS consists of a specialized transmission line known as a dielectric waveguide. Similar to leaky-wave antennas, electromagnetic waves within the waveguide are emitted by pinching small separate dielectric particles at specific points along its length. These dielectrics are typically attached to the tips of plastic clips or clothespins, forming structures known as pinching antennas (PAs). By leveraging the waveguide’s ability to transmit or receive radio waves at any pinched point, PASS enables wireless communication precisely where needed [8].
The waveguide can be pinched at multiple points to provide coverage in specific locations, allowing for flexible and dynamic antenna deployment. Signal emission at a particular location ceases immediately upon removing the pinching antenna, making the process as simple as adding or removing a clothespin from a clothesline, as illustrated in Figure 1. This adaptability enables PASS to establish communication zones tailored to users’ locations and environmental conditions, offering exceptional flexibility in antenna deployment [8].
Unlike conventional flexible-antenna systems, the length of a PASS waveguide can be extended arbitrarily, allowing PAs to be positioned closer to users to establish strong line-of-sight (LoS) links [9]. This adaptability effectively creates a near-wired connection, ensuring favorable channel conditions and shifting the focus of wireless communication networks from the “last mile” to the more manageable “last meter”. This unique capability is reflected in the name “PASS”, highlighting its ability to provide the shortest and most efficient wireless connections between the PAs and the wireless transceivers, which effectively pass wireless signals with minimal loss.
@@ I-B Key Advantages for Employing PASS in Wireless Communications
  • Enhanced LoS Stability and “Last Meter” Communications: By extending the waveguide and enabling the flexible deployment of PAs, PASS effectively mitigate large-scale path loss and establish stable LoS communication environments. This capability enhances “last meter” communications and expands network coverage, which is not achievable with other flexible-antenna technologies.
  • Flexible Pinching Beamforming Gain: PASS control signals by optimizing the locations and the numbers of PAs, which enables highly flexible and scalable pinching beamforming. This capability allows for precise management of signal strength and interference, effectively meeting stringent service requirements.
  • Utilization of Near-Field Benefits: The extended aperture of the waveguide places users within the near-field region of PASS, which enables the use of near-field electromagnetic properties, such as beam focusing, to harvest the spatial degrees of freedom and enhance wireless transmission efficiency [10].
  • Cost-Effective and Scalable Implementation: PASS are simple to implement and economical to deploy, not only because of the low cost of waveguides, but also because PASS only require adding or removing dielectric materials at selected points along the waveguide [8], which makes PASS a scalable choice for practical network design.
These advantages have fueled growing research interest in PASS. The performance of PASS has been analyzed in terms of its ergodic rate [9] and its array gain [11]. Furthermore, various activation algorithms have been proposed to optimize the locations of PAs along a single waveguide; see [9] and relevant references. However, despite these initial efforts, the design of PASS for wireless communications remains in its early stages. This gap motivates us to provide a systematic introduction to PASS. In this article, we first discuss how PASS differ from conventional multi-antenna systems. We then propose several novel beamforming architectures to facilitate the design of pinching beamforming. This is followed by an exploration of promising application scenarios for PASS and an evaluation of their performance. Finally, we highlight the open challenges of PASS to inspire future research in this promising research area.
@@ II-A Physical Principle
Unlike conventional electronic antennas which rely on electronic devices to convert alternating electric currents into electromagnetic (EM) waves, the physical principle of PAs is based on the phenomenon of EM coupling that arises when a separate dielectric is brought into close proximity with a dielectric waveguide [8]. This coupling enables the power exchange between the waveguide and the separate dielectric, as illustrated in the bottom-right of Fig. 1. Therefore, by pinching a separate dielectric onto the waveguide, the EM waves originally confined within the waveguide are coupled into the separate dielectric, where they are subsequently radiated into free space, thereby forming an “antenna”. Unlike conventional electronic antennas, PAs are entirely passive, simpler to manufacture and deploy, and hence provide greater position flexibility.
By modeling the pinching antenna as an open-ended directional coupler, the power exchange caused by the EM coupling can be characterized by the coupled-mode theory. Let $P_{\mathrm{guide}}$ and $P_{\mathrm{pinch}}$ denote the normalized power in the waveguide and the pinched separate dielectric, respectively. Their relationship is approximately given by [12]
@@
where $L$ is the coupling length shown in Fig. 1, $\kappa$ is the coupling coefficient, and $F\leq 1$ is the maximum coupling efficiency. The factors $\kappa$ and $F$ are influenced by various factors, including the effective refractive index and the transverse modes of both the waveguide and the separate dielectric. It can be observed from the above formulas that the power emitted by a PA can be controlled by adjusting the coupling length $L$. When the waveguide and the separate dielectric share the same effective refractive index, the coupling efficiency $F$ reaches its maximum value of $1$, making it possible to radiate full power from a single PA with a coupling length of $\pi/(2\kappa)$.
@@ II-B Signal Model
As shown in Fig. 1, the composite channel from the base station (BS) to the user in the PASS system consists of two components: in-waveguide propagation and free-space propagation. Compared to free-space propagation, in-waveguide propagation incurs almost negligible path loss, e.g., approximately 0.01-0.03 dB/m for a circular copper waveguide at 15 GHz [13]. Consequently, when multiple PAs are deployed on a waveguide, the signal received at the user from the $n$-th PA through the LoS channel can be expressed mathematically as:
@@
Based on the physical principle of PAs, the power model of PASS, signified by $P_{n}$, is significantly different from that of conventional multi-antenna systems. Specifically, with conventional antennas, the power of the signal emitted from each antenna is actively controlled by electronic devices including radio-frequency (RF) chains and power splitters. In contrast, for PAs, the transmit power is passively determined by the coupling length with the waveguide. Furthermore, when multiple PAs are pinched along a waveguide, the power radiated by subsequent antennas is influenced by the power radiated by preceding ones, leading to a coupled power model. While the coupling length of a PA is not as easily adjustable as the electronic parameters of conventional antennas since PAs are typically passive, it is still possible to design and manufacture PAs with specific coupling lengths to achieve a desired power model. In the following, we introduce two common power models for PAs.
  • Equal Power Model: In this model, the PAs on the waveguide radiates an equal amount of power. This can be achieved by designing the PAs with varying coupling lengths, such that the earlier PAs radiate a smaller proportion of the power in the waveguide, while the later PAs radiate a larger proportion of the remaining power.
  • Proportional Power Model: In this model, each PA is designed with an identical coupling length, thereby radiating the same fixed proportion of the power left in the waveguide. As a result, the power radiated by each PA decreases progressively along the waveguide, where the power radiated by later PAs is proportional to that of earlier ones by a fixed ratio.
These two power models offer different pros and cons. Compared to the proportional power model, where the latter PAs may radiate only a smaller amount of power, the equal power model ensures that all PAs are effectively utilized. The equal power model also simplifies the performance analysis of the PASS system due to its simple mathematical formulation, which is useful to identify the fundamental limits of PASS, and hence has contributed to its widespread adoption in the literature. In contrast, the proportional power model requires less stringent manufacturing specifications, as all PAs share the same design, thus reducing production complexity and costs.
@@ II-C Communication System Design
From a communication system design perspective, it is crucial to position the PAs at suitable locations to maximize communication performance. On the one hand, the PAs should be deployed close to the user to mitigate the significant signal attenuation caused by free-space path loss and blockage. On the other hand, when multiple PAs are deployed along the waveguide, they form an antenna array, enabling pinching beamforming to enhance signal strength and reduce interference. By fine-tuning the positions of PAs along the waveguide, the phase shifts of the signals, due to the in-waveguide and free-space propagation, can be adjusted, thereby optimizing the beamforming gain. Leveraging the flexibility and portability of PAs, their positions can be determined either actively by the BS or passively by the user, leading to the following two designs:
  • BS-centric Design: In this design, the positions of PAs are determined by the BS to optimize specific performance metrics of the overall PASS system, such as spectral efficiency and user fairness. In dynamic scenarios, the positions of PAs can be adjusted actively based on the locations of all users in two ways: either by moving them along a pre-installed track parallel to the waveguide for continuous activation, or by selectively activating PAs at the pre-configured locations on the waveguide for discrete activation.
  • User-centric Design: In this design, users can pinch or release portable PAs onto the waveguide as needed to enhance communication service in their local area. Whenever a PA is pinched or released by a user, the baseband processing strategy needs to be optimized and updated based on the new propagation environment to avoid the performance degradation to other users.
@@ III-A Non-Multiplexing Architecture
A schematic illustration of the non-multiplexing architecture is shown in Fig. 2. In this architecture, the data streams to be transmitted are not multiplexed at the baseband, meaning that each waveguide is fed individually by a single data stream. In this setup, the primary role of the baseband processing is to manage the power allocation across these data streams to optimize transmission efficiency. The key benefit of this design is to reduces implementation complexity. Compared to the conventional multi-antenna systems, the unique capability of pinching beamforming, achieved by adjusting the positions of PAs, can be utilized to enhance communication performance. This non-multiplexing architecture is particularly well-suited for those commonly encountered scenarios, where each waveguide serves geographically isolated areas, minimizing the need for advanced interference management between waveguides.
@@ III-B Multiplexing Architecture
As shown at the bottom of Fig. 3, the multiplexing architecture multiplexes multiple data streams before feeding them into the waveguides, thus facilitating the joint transmit and pinching beamforming. This architecture can be seen as an extension of the conventional MIMO transmission architecture, where each electronic antenna is replaced by a waveguide. The key difference between the two architectures lies in the point of signal radiation. In conventional MIMO systems, signals are radiated directly from the electronic antennas into free space. In contrast, in the multiplexing architecture of PASS, signals are radiated at specific locations along the waveguide via PAs. Consequently, unlike conventional MIMO, which relies on a massive number of electronic antennas to achieve high spatial degrees of freedom (DoFs) [14], the multiplexing architecture of PASS significantly reduces the requirement for numerous waveguides because a large number of low-cost PAs can be easily installed on a single waveguide. Based on the above discussion, the multiplexing architecture for PASS can be designed in analogy to conventional MIMO architecture but with distinct differences in implementation and design focus. In the following, we introduce three basic multiplexing architectures, as shown in Fig. 3, namely sub-connected architecture, fully-connected architecture, and phase shifter (PS)-based fully-connected architecture.
  • Sub-connected Architecture: This is the simplest multiplexing architecture, where each waveguide is connected to a dedicated RF chain and carries signals that are multiplexed and precoded by baseband digital beamforming. While this architecture is analogous to the fully-digital architecture used in conventional MIMO systems, it is referred to as “sub-connected” because the signals from each RF chain are emitted exclusively by those PAs deployed on the corresponding waveguide.
  • Fully-connected Architecture: In this architecture, the signals from each RF chain are fed into all waveguides with the aid of power splitters. The fully-connected architecture addresses two key limitations of the sub-connected design. First, it enables each RF chain to connect to all PAs across all waveguides, thereby fully using the overall spatial DoFs to enhance communication performance. Second, the number of RF chains no longer needs to match the number of waveguides. This flexibility allows for either reducing the number of RF chains to save energy, as hybrid beamforing in millimeter-wave networks, or increasing it to further exploit the additional spatial DoFs provided by the massive PAs.
@@
IV “PASS” Signals for 6G and Beyond